### v1

@@ 6.5 Vulnerability Ranking
MISTRAL-7B
1 $\mathcal{DF}$ Dereference failure: NULL pointer 6,318 33.19%
2 $\mathcal{BO}$ Buffer overflow on scanf 5,169 27.15%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 2,474 13.00%
4 $\mathcal{DF}$ Dereference failure: array bounds violated 747 3.92%
5 $\mathcal{DF}$ Array bounds violated: lower bound 621 3.26%
6 $\mathcal{AO}$ Arithmetic overflow on sub 475 2.50%
7 $\mathcal{DF}$ Array bounds violated: upper bound 452 2.37%
8 $\mathcal{DF}$ Dereference failure: forgotten memory 413 2.17%
9 $\mathcal{AO}$ Arithmetic overflow on add 402 2.11%
10 $\mathcal{BO}$ Buffer overflow on sscanf 393 2.06%
GPT-4
1 $\mathcal{BO}$ Buffer overflow on scanf 600 34.70%
2 $\mathcal{DF}$ Dereference failure: NULL pointer 493 28.51%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 210 12.15%
4 $\mathcal{AO}$ Arithmetic overflow on sub 61 3.53%
5 $\mathcal{AO}$ Arithmetic overflow on floating-point ieee_mul 57 3.30%
6 $\mathcal{AO}$ Arithmetic overflow on add 56 3.24%
7 $\mathcal{DF}$ Dereference failure: forgotten memory 55 3.18%
8 $\mathcal{DF}$ Dereference failure: array bounds violated 51 2.95%
9 $\mathcal{DF}$ Array bounds violated: lower bound 34 1.97%
10 $\mathcal{AO}$ Arithmetic overflow on mul 25 1.45%
While dereference failures and buffer overflows pose risks, their severity and frequency vary significantly. For instance, GEMMA-7B shows a disproportionately high incidence of NULL pointer dereference failures at $65.76\%$, suggesting specific weaknesses in its memory management capabilities. Arithmetic overflows appear consistently for all models, mostly populating the bottom 5 ranks. They vary in terms of the specific operations (add, sub, mul) affected, indicating differing arithmetic handling across the models. Interestingly, lallama2-13B stands out as the only model below $10\%$ on scanf() related violations, with Gemini-pro following suit around $11\%$. However, similar to GEMMA-7B, both models exhibit a disproportionately high incidence of NULL pointer dereference failures.
The consistent appearance of certain errors across different models emphasizes the need for comprehensive testing and validation frameworks to address these recurrent issues before deployment. While all models exhibit the same vulnerabilities, significant differences in the frequency and type of other vulnerabilities, such as arithmetic overflows, suggest that model-specific optimizations and enhancements are necessary.
6.6 LLM Ranking: Which model is the most secure
To compare which model is the “worst” or the “best” when it comes to secure coding—and to do this as fairly as possible—we will investigate several metrics, such as the ratio of verification results, average property violation per file, and average property violation per line of code.
@@ 6.6 LLM Ranking: Which model is the most secure
GPT4 0.0165 3 11.40% 2 50.80% 36.50% 3.40
Llama2-13B 0.0234 2 13.36% 1 47.50% 35.58% 3.62
Mistral_7B 0.0254 7 7.11% 4 62.00% 28.19% 3.07
Codellama_13B 0.0260 1 15.24% 3 52.34% 30.12% 4.13
Falcon180B 0.0291 8 6.49% 5 62.10% 28.61% 3.38
GPT35 0.0295 6 7.57% 7 64.25% 26.65% 4.42
Gemini_Pro 0.0305 5 9.56% 6 63.63% 24.39% 4.70
Gemma7B 0.0437 4 11.29% 8 69.29% 15.28% 4.20
@@
$\boldsymbol{\mathcal{VS}}$: Verification Success; $\boldsymbol{\mathcal{VF}}$: Verification Failed; $\boldsymbol{\mathcal{VU}}$: Verification Unknown (Timeout)
The results indicate that there is no clear winner. Mistral-7B, despite having the fewest property violations per file, writes shorter code, reducing its likelihood of coding errors. However, this model also performs poorly in the $\mathcal{VS}$ metric, with only 7.11% of its samples categorized as being free of vulnerabilities. Codellama-13B achieved the highest $\mathcal{VS}$ rate, followed by Llama2-13B, and their $\mathcal{VF}$ ratio is ranking is third and second respectively, which is a good result for the LLama family. Still, it is best to remember that nearly half of their samples had vulnerabilities. Moreover, their $\mathcal{VU}$ is fairly high at $30\%$ and $35\%$, which means that with further verification, there is still a chance that other models will take the lead.
GPT-4 outperforms GPT-3.5 while showing the highest $\mathcal{VU}$ percentage under current ESBMC settings, indicating its ability to produce more complex and longer outputs. It is important to note that this complexity is not reflected by the CC number as discussed earlier, which confirms the criticism towards Cyclomatic Complexity by practitioners. While GPT-4 ranks third in $\mathcal{VS}$ and second in $\mathcal{VF}$, it finishes first with an average property violation per line. This might be the fairest way to compare models, as the more lines, the more chances to have vulnerabilities, while this metric doesn’t punish models producing shorter codes.
There is no definitive winner in this analysis; however, Gemma-7B, Gemini-Pro, and GPT-3.5—with the current verification settings—has the highest $\mathcal{VF}$ ratios and highest average property violation both per line and file. This unveils an interesting and unexpected finding: empirically, a slight correlation between having high cyclomatic complexity and vulnerable coding patterns can be observed. As Table 10 shows, these three models had the worst CC statistics. The correlation is, however, not clear, and there are certainly other factors. For example, CodeLlama-13B, with a maximum CC number of 94, did not show poor performance in the verification results.
It is important to underline that it might be tempting to speculate on a winner. However, having such a high verification failed ratio is unacceptable from a SE perspective for any model. All models surpassed the $\mathcal{VF}$ threshold of 47%, indicating that nearly half or more of the generated programs are vulnerable. The conclusions of this analysis must be clear: Using code generated by the state-of-the-art Large Language Models, without any additional framework for validation and vulnerability analysis, carries severe risks. While LLMs can be useful for automating simple tasks and scripting, directly including such codes in production software without oversight from experienced software engineers is irresponsible and should be avoided.
@@ 7.1 Future Research Directions
The dataset containing all $265,000$ C program files, along with the .json and .csv files are published on GitHub [^6] https://github.com/FormAI-Dataset. The dataset is specifically prepared to support Machine Learning and fine-tuning. The absence of false positives makes the dataset suitable for benchmarking the effectiveness of various static and dynamic analysis tools and ML-based classifiers. Training on data that is void of false negatives is really important, so we indicate in the .json file for each file whether the verification process has finished, as such files do not contain the vulnerabilities detectable by ESBMC. The diverse structure of the C programs generated in the FormAI-v2 dataset made it excellent for an unexpected use case: fuzzing different applications. We uncovered and reported over thirteen bugs while executing ESBMC using the FormAI dataset. After validating these issues, ESBMC developers managed to resolve them. These included errors in the goto-cc conversion and the creation of invalid SMT solver formulas. Additionally, we identified bugs in the CBMC [111] and the Clang compiler, which failed to compile several programs, while GNU C had no issue. We promptly communicated these findings to the respective developers. The FormAI-v2 dataset aims to be a useful resource for training machine learning algorithms to possess the capabilities of the ESBMC module.
@@
  • It would be important to investigate why programs under “Verification Successful” are void of vulnerabilities. Is it because of better coding practices or simply because for example they don’t take user input, thereby avoiding buffer overflows?
@@
  • As our motivation section showcased, GPT-4 did not excel at avoiding and fixing the vulnerability in the example. How do different LLMs compare in understanding, correctly fixing, and detecting coding errors?
  • We aim to further grow the FormAI dataset including more state-of-the-art models, and also by increasing the number of samples for each LLM to have an overall larger dataset.
@@
While we can partially address the last question, noting the use of insecure functions and poor input sanitization in handling user inputs, exploring this issue across various domains, such as networking or cryptography would be beneficial.
@@ 7.2 Limitations and Threats to Validity
With a larger timeout setting, ESBMC might find slightly more vulnerabilities in a given program. Whether the verifier can finish the process under a given timeout is up to the available computational capacity. The same parameter setting can yield a higher or lower detection rate on different architectures. To find all errors detectable by ESBMC, unwind must be set to infinite, and ESMBC must complete the verification process. As we provided the original C programs and the instructions on how to run ESBMC, researchers who invest additional computational resources have the potential to enhance our findings. As the “Verification Unknown” category still contains samples for every model, the current results are strongly bound to the percentage of vulnerable files LLMs produce.
While ESBMC is a robust tool for detecting many types of errors in C, it is not currently suited to detect design flaws, semantic errors, or performance issues. As such, more vulnerabilities might be present besides the detected ones in the code. Thus we recommend that the training and fine-tuning applications to be restricted to the vulnerabilities detectable by ESBMC on this dataset.
@@ 8 Conclusions
In this research, we analyzed eight state-of-the-art Large Language Models to assess their likelihood of introducing vulnerabilities during neutral prompt based code generation. The models included in our analysis were Mistral-7B, Falcon-180B, GPT-4, Llama2-13B, Codellama-13B, Gemma-7B, GPT-3.5, and Gemini-Pro. We employed a zero-shot prompting method to encompass numerous programming scenarios for C code generation. These programs constitute the FormAI-v2 dataset, containing $256,000$ independent compilable C programs.
@@
We used the Efficient SMT-based Bounded Model Checker (ESBMC), a state-of-the-art formal verification tool, to identify vulnerabilities. Each program was given a verification period of 500 seconds with the unwinding parameter set to infinite, uncovering a total of 684,227 vulnerabilities. Overall $67\%$ of the codes were vulnerable. Detailed labelling of each sample—including filename, type of vulnerability, function name, error type, and source code—is documented in a .json file, as detailed in Appendix Fig. 11, to facilitate the dataset’s use in machine learning applications.
@@
Additionally, the FormAI-v2 dataset proved instrumental for fuzzing various applications, leading to the identification of multiple bugs in ESBMC, CBMC, and the Clang compiler. We have identified 42 distinct CWE identifiers, six of which are featured on MITRE’s Top 25 list for 2023. Notably, ”CWE-787 Out-of-bounds Write,” the top vulnerability on the list, was confirmed in at least 181,122 cases. These findings provide clear answers to our research questions:
@@
While the literature reveals significant variations in the ability of these models to solve tasks, this is not mirrored in their susceptibility to produce vulnerabilities in source code. Our findings conclusively show that despite differences among the examined models, in terms of generating code, they all consistently introduce severe vulnerabilities when prompted with simple coding tasks. Our study indicates that despite the impressive capabilities of Large Language Models in code generation, employing their output in production requires meticulous risk assessment. Relying on these models without expert oversight in a production context is inadvisable.
Appendix
Switch Description
–timeout <seconds> This switch establishes a time limit in seconds for the analysis. If the verification process exceeds this duration, ESBMC will automatically halt the analysis.
–k-induction <number> This enables k-induction proof, a formal verification technique to prove properties of loops.
–unlimited-k-steps Allows the k-induction process to proceed without a predefined limit on the number of iterations, aiming for more thorough verification.
–unwind <number> Sets the unwind limit for loops and recursion. This specifies how many times loops or recursive functions should be unwound during the analysis.
–no-unwinding-assertions Disables assertions that check whether the loop or recursion unwinding was sufficient, useful when the unwind bound is considered adequate.
–falsification ESBMC prioritizes finding violations of the properties, and the tool terminates as soon as it discovers a counterexample.
–overflow Enables checking for arithmetic overflows within the program, helping detect when operations exceed the data type limits.
–memory-leak-check Activates the detection of memory leaks, ensuring that all allocated memory is properly freed before program termination.
–multi-property Enables the verification of multiple properties simultaneously, optimizing the analysis by checking all specified properties in a single run.
–show-stacktrace When an error occurs, this option provides a stack trace, useful for debugging by showing the sequence of function calls leading to the error.
–verbosity 6 Sets the verbosity level of the output to detailed, offering a deeper insight into the tool’s operations and the verification process.
–incremental-bmc Enhances standard BMC by gradually increasing the bound depth, beginning at a low level and continuing until a bug is detected, a specified depth is reached, or resources are exhausted.
Table 16: Description of Investigated ESBMC Switches.
Figure 11: Example from the JSON file, demonstrating the labelling process. The source code object has been omitted.
@@
In this study, a total of $265,000$ C samples were generated and examined. The findings and all the generated C samples are available for access and download from the project’s website at https://github.com/FormAI-Dataset.

### v2

@@ 6.3 Vulnerability Ranking
Falcon2-11B
1 $\mathcal{DF}$ Dereference failure: NULL pointer 12 8631286312\,86312 863 40.71%
2 $\mathcal{BO}$ Buffer overflow on scanf 6 59365936\,5936 593 20.87%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 4 51545154\,5154 515 14.29%
4 $\mathcal{DF}$ Dereference failure: array bounds violated 1 26612661\,2661 266 4.01%
5 $\mathcal{DF}$ Dereference failure: forgotten memory 1 10611061\,1061 106 3.50%
6 $\mathcal{DF}$ Array bounds violated: upper bound 1 07410741\,0741 074 3.40%
7 $\mathcal{AO}$ Arithmetic overflow on add $762$ 2.41%
8 $\mathcal{DF}$ Array bounds violated: lower bound $613$ 1.94%
9 $\mathcal{AO}$ Arithmetic overflow on sub $561$ 1.78%
10 $\mathcal{AO}$ Arithmetic overflow on mul $459$ 1.45%
Dereference failures, particularly NULL pointer dereferences, are among the most prevalent vulnerabilities across all LLMs. This is due in part to the varied and often unsafe examples of pointer usage in training datasets, combined with the inherent complexity of dynamic memory management in C. LLMs rely on pattern recognition rather than deep understanding, which leads them to mishandle pointers and fail to replicate the nuanced behavior of real-world applications. This results in frequent dereference issues and flawed pointer handling, highlighting significant risks when deploying LLM-generated code in critical systems where security and reliability are paramount.
The severity and frequency of these vulnerabilities vary significantly among models. For instance, Gemma-7B exhibits a notably high rate of NULL pointer dereference failures at 60.50%, indicating substantial weaknesses in memory management. Arithmetic overflows also consistently appear across all models in the top 10 list, and differ based on specific operations (addition, subtraction, multiplication), underscoring varied arithmetic handling. Notably, Llama2-13B stands out with less than 10% of scanf violations, with Gemini Pro 1.0 close behind at approximately 11%; however, both models, like Gemma-7B, show high rates of NULL pointer dereference failures.
The consistent occurrence of certain errors across different models underscores the need for comprehensive testing and validation frameworks to address these recurring issues before deployment. While all models share similar vulnerabilities, significant differences in the frequency and types of other vulnerabilities—such as arithmetic overflows—suggest that model-specific optimizations and enhancements are necessary. To mitigate these risks, developing enhanced training methodologies focused on robust memory handling is crucial. Implementing advanced code analysis tools and frameworks is also essential to detect and rectify vulnerabilities before deployment for real-world applications.
6.4 LLM Ranking: Which model is the most secure coder
To compare which model is performing the “worst” or the “best” when it comes to secure coding—and to do this as fairly as possible—we will investigate several metrics, such as the ratio of verification results, average property violation per file, and average property violation per line of code.
@@ 6.4 LLM Ranking: Which model is the most secure coder
GPT-4o-mini 0.0165 3 4.23% 2 57.14% 36.77% 3.40
Llama2-13B 0.0234 2 12.36% 1 51.30% 31.78% 3.62
Mistral-7B 0.0254 7 8.36% 4 62.08% 25.88% 3.07
CodeLlama-13B 0.0260 1 15.48% 3 52.71% 29.52% 4.13
Falcon-180B 0.0291 8 6.48% 5 62.07% 28.67% 3.38
GPT-3.5-turbo 0.0295 6 7.29% 7 65.07% 26.09% 4.42
Gemini Pro 1.0 0.0305 5 9.49% 6 63.91% 24.13% 4.70
Gemma-7B 0.0437 4 11.62% 8 67.01% 16.30% 4.20
@@
$\boldsymbol{\mathcal{VS}}$: Verification Success; $\boldsymbol{\mathcal{VF}}$: Verification Failed; $\boldsymbol{\mathcal{VU}}$: Verification Unknown (Timeout).
Best performance in a category is highlighted with bold and/or Rank.
The results indicate that there is no clear winner. Mistral-7B, despite having the fewest property violations per file, writes shorter code, reducing its likelihood of coding errors. However, this model also performs poorly in the $\mathcal{VS}$ metric, with only 8.36% of its samples categorized as being free of vulnerabilities. CodeLlama-13B achieved the highest $\mathcal{VS}$ rate, followed by Llama2-13B, and their $\mathcal{VF}$ ratio ranking is third and second respectively, which is a good result for the Llama family. Still, it is best to remember that nearly half of their samples had vulnerabilities. Moreover, their $\mathcal{VU}$ is fairly high at $30\%$ and $32\%$, which means that with further verification, there is still a chance that other models will take the lead.
GPT-4o-mini outperforms GPT-3.5-turbo while showing the highest $\mathcal{VU}$ percentage under current ESBMC settings, indicating its ability to produce more complex and longer outputs. It is important to note that this complexity is not reflected by the CC number as discussed earlier, which confirms the criticism towards Cyclomatic Complexity by practitioners. While GPT-4o-mini ranks third in $\mathcal{VS}$ and second in $\mathcal{VF}$, it finishes first with an average property violation per line. This might be the fairest way to compare models, as the more lines, the more chances to have vulnerabilities, while this metric doesn’t punish models producing shorter codes. While there is no definitive winner in this analysis, Gemma-7B, Gemini-Pro, and GPT-3.5-turbo—with the current verification settings— have the highest $\mathcal{VF}$ ratios and highest average property violation both per line and file which indicates that these models are performing worse in our test.
It is important to underline that it might be tempting to speculate on a winner, having such a high verification failed ratio is unacceptable from an SE perspective for any model. All models surpassed the $\mathcal{VF}$ threshold of 50%, indicating that nearly half or more of the generated programs are vulnerable. The conclusions of this analysis must be clear: Using code generated by the state-of-the-art Large Language Models, without any additional framework for validation and vulnerability analysis, carries severe risks. While LLMs can be useful for automating simple tasks and scripting, directly including such codes in production software without oversight from experienced software engineers is irresponsible and should be avoided.
@@ 7.1 Future Research Directions
The dataset, consisting of 331 000331000331\,000331 000 C program files and their corresponding vulnerability classifications, is available on GitHub [^6] https://github.com/FormAI-Dataset. The dataset is well-suited for machine learning applications and fine-tuning LLMs due to its large size. Moreover, the diverse structure of the C programs generated by various LLMs in the FormAI-v2 dataset makes it ideal for an unexpected use case: fuzzing different applications. We discovered and reported seventeen bugs in the ESBMC application, including issues related to unsigned overflow checks, SMT solver problems, conversion errors in the GOTO converter, and flaws in implementing the k-induction proof rule. Furthermore, we identified bugs in the CBMC [116] tool while using the FormAI-v2 dataset and promptly communicated these findings to the respective developers. After validating the reported issues, the ESBMC developers have already resolved thirteen.
@@
  • It would be important to investigate why programs under “Verification Successful” are void of vulnerabilities. Is it because of better coding practices or simply because, for example, they don’t take user input, thereby avoiding buffer overflows?
@@
  • As our motivation section showcased, GPT-4o-mini did not excel at avoiding and fixing the vulnerability in the example. How do different LLMs compare in understanding, correctly fixing, and detecting coding errors?
  • We aim further to grow the FormAI dataset, including more state-of-the-art models, and increase the number of samples for each LLM to have an overall larger dataset.
@@
While we can partially address the last question, noting the use of insecure functions and poor input sanitation in handling user inputs, exploring this issue across various domains, such as networking or cryptography, would be beneficial.
@@ 7.2 Limitations and Threats to Validity
ESBMC might find slightly more vulnerabilities in a given program with a larger timeout setting. Whether the verifier can finish the process under a given timeout is up to the available computational capacity. The same parameter setting can yield a higher or lower detection rate on different architectures. To find all errors detectable by ESBMC, unwind must be set to infinite, and ESMBC must complete the verification process. As we provided the original C programs and the instructions on how to run ESBMC, researchers who invest additional computational resources have the potential to enhance our findings. As the “Verification Unknown” category still contains samples for every model, the current results are strongly bound to the percentage of vulnerable files LLMs produce.
While ESBMC is a robust tool for detecting many types of errors in C, it is not currently suited to detect design flaws, semantic errors, or performance issues. As such, more vulnerabilities might be present besides the ones detected in the code. Thus, we recommend that the training and fine-tuning applications be restricted to the vulnerabilities detectable by ESBMC on this dataset.
All programs shorter than 50 lines were removed from the dataset. However, researchers interested in smaller programs can still find all programs under 50 lines, generated by GPT-3.5-turbo, in the original FormAI-v1 dataset.
@@ 8 Conclusions
This research analyzed nine state-of-the-art Large Language Models to assess their likelihood of introducing vulnerabilities during neutral prompt-based code generation, and to compare their performance. The models included in our analysis were Mistral-7B, Falcon-180B, Falcon2-11B GPT-4o-mini, Llama2-13B, CodeLlama-13B, Gemma-7B, GPT-3.5-turbo, and Gemini-Pro. We employed a zero-shot prompting method to encompass numerous programming scenarios for C code generation. These programs constitute the FormAI-v2 dataset, containing 331 000331000331\,000331 000 independent compilable C programs.
@@
We used the Efficient SMT-based Bounded Model Checker (ESBMC), a state-of-the-art formal verification tool, to identify vulnerabilities. Each program was given a verification period of 500 seconds with the unwinding parameter set to infinite, uncovering a total of 765 366765366765\,366765 366 vulnerabilities. Overall $62.07\%$ of the codes were vulnerable. Detailed labeling of each sample—including filename, type of vulnerability, function name, error type, and source code—is documented in a .json file, as detailed in Appendix Fig. 9, to facilitate the dataset’s use in machine learning applications.
@@
Additionally, the FormAI-v2 dataset proved instrumental for fuzzing various applications and identifying multiple bugs in ESBMC and CBMC. These findings provide clear answers to our research questions:
RQ1: How does the security of LLM-generated code differ across various models? Answer: CodeLlama-13B, Llama-13B, and GPT-4o-mini perform slightly better, but all examined models notoriously introduce vulnerabilities into the C code they generate at unacceptable rates. Our research revealed that all examined models introduced vulnerabilities in at least 50% of the generated code.
RQ2: What are the most typical vulnerabilities introduced by different LLMs during code generation (focusing on C)? Answer: Dereference failures and buffer overflow issues are the most prevalent vulnerabilities across all models, ranking arithmetic overflow as the third most common type. No model is completely free from any of the examined vulnerabilities; the variations lie in the frequency of occurrence.
@@
While the literature reveals significant variations in these models’ ability to solve tasks, this is not mirrored in their susceptibility to produce vulnerabilities in source code. Our findings conclusively show that despite differences among the examined models in terms of generating code, they all consistently introduce severe vulnerabilities when prompted with simple coding tasks. Our study indicates that despite the impressive capabilities of Large Language Models in code generation, employing their output in production requires detailed risk assessment. Relying on these models without expert oversight in a production context is inadvisable.
Acknowledgement
We extend our sincere thanks to the anonymous reviewers for their valuable feedback, which has significantly improved the quality of this paper. This research is supported by the Technology Innovation Institute (TII), Abu Dhabi. Additionally, partial support is provided by the EPSRC grant EP/T026995/1, titled “EnnCore: End-to-End Conceptual Guarding of Neural Architectures” under the Security for All in an AI-enabled Society initiative. This work is also partially supported by the TKP2021-NVA Funding Scheme under Project TKP2021-NVA-29.
@@ Acknowledgement
This study generated and examined a total of 331 000331000331\,000331 000 C samples. The findings and all the generated C samples are available for access and download from the project’s website at https://github.com/FormAI-Dataset.
@@
Appendix
Figure 9: Example JSON Labels for a GPT-3.5-turbo Generated Sample: FormAI-v2 dataset